You are building a rectangular deck. The area of the deck should be 250 ft squared. You want the length of the deck to be 5 ft longer than twice its width. What should the dimensions of the deck be?
Any help is appreciated

Answer by rothauserc(4718)

You are building a rectangular deck. The area of the deck should be 250 ft squared. You want the length of the deck to be 5 ft longer than twice its width.
let w be the width and 2w+5 is the length, then
w(2w+5) = 250
2w^2+5w-250=0
(2w+25) * (w-10) = 0 and
w=10 and l = 25